I have been troubleing over this Question for awhile now tried solving it with question No. 63487 from this website but the numbers were close but not exact.
3^x=7.5^(x-2)

Answer by ankor@dixie-net.com(12692)

3^x = 7.5^(x-2)
:
Write it using the log (base 10) equiv of exponents:
x * log(3) = (x-2)* log(7.5)
:
Find the logs of 3 and 7.5:
.47712x = .87506(x-2)
:
Mult what's inside the brackets:
.47712x = .87506x - 1.75012
:
Subtract .87506x from both sides
.47712x - .87506x = - 1.75012
-.39794x = -1.75012
x = -1.75012/-.39794
x = + 4.3979
:
Check using a calc:
3^4.3979 = 125.4
7.5^2.3979 = 125.4 also
:
Often you won't come out exactly, but if you go to 4 decimal places it will be very close.